# Audio Personas: Augmenting Social Perception via Body-Anchored Audio Cues

### v1

@@ 7. Conclusion
We introduced Audio Personas, a novel concept to augment face-to-face interactions with body-anchored audio cues. Through a formative study with designers, we uncovered potential use cases and elicited design patterns. Our in-lab study further revealed that audio personas can influence how people formed social impressions. Individuals with a positive audio persona were deemed as more socially attractive, likable, and less threatening than ones with a negative audio persona. Audio personas open new possibilities for engaging dynamic multisensory channels to enhance face-to-face social interactions and this paper laid out key building blocks for future explorations.
@@ Acknowledgements.
We would like to thank Keshav Rastogi, Muhammad Ali, and Brian Beams for their support on the user study logistics. We thank Anna Matsumoto and Emmanuel Corona-Moreno for their help on the video production. We would also like to thank everyone who provided feedback on the paper draft: Danyang Fan, Han Guo, Eugy Han, Saehui Hwang, Monique Santoso, Tara Srirangarajan, Olivia Tomassetti, Ahad Rauf, Savannah Cofer, Portia Wang, Yifei Cheng, Yiran Zhao, and Ran Zhou. The lead author of this work is supported by the Knight-Hennessy Scholarship.References
@@
- Goffman (2023) Erving Goffman. 2023. The presentation of self in everyday life. In Social theory re-wired. Routledge, 450–459.
@@
- Oxford English Dictionary (2023) Oxford English Dictionary. 2023. persona (n.), sense 2.a. https://doi.org/10.1093/OED/1145536616

### v2

@@ 7. Conclusion
We introduced Audio Personas, a novel concept to augment face-to-face interactions with body-anchored audio cues. Our in-lab study revealed that audio personas can influence how people formed social impressions. Individuals with a positive audio persona were deemed as more socially attractive, likable, and less threatening than ones with a negative audio persona. Through a design study with audio designers, we further uncovered potential use cases and elicited design patterns. Audio personas open new possibilities for engaging dynamic multisensory channels to enhance face-to-face social interactions and this paper laid out key building blocks for future explorations.
@@ Acknowledgements.
We would like to thank Keshav Rastogi, Muhammad Ali, and Brian Beams for their support on the user study logistics. We thank Anna Matsumoto and Emmanuel Corona-Moreno for their help on the paper video. We would also like to thank everyone who provided feedback on this project: Ethan Buck, Prof. Ana Tajadura-Jiménez, Danyang Fan, Han Guo, Eugy Han, Saehui Hwang, Monique Santoso, Tara Srirangarajan, Olivia Tomassetti, Ahad Rauf, Savannah Cofer, Portia Wang, Yifei Cheng, Yiran Zhao, and Ran Zhou. The lead author of this work is supported by the Knight-Hennessy Scholarship.References
@@
- Frid et al. (2016) Emma Frid, Roberto Bresin, Paolo Alborno, and Ludvig Elblaus. 2016. Interactive sonification of spontaneous movement of children—cross-modal mapping and the perception of body movement qualities through sound. Frontiers in neuroscience 10 (2016), 521.
@@
- Hall (1963) Edward T Hall. 1963. A system for the notation of proxemic behavior. American anthropologist 65, 5 (1963), 1003–1026.
@@
- Hayduk (1983) Leslie A Hayduk. 1983. Personal space: Where we now stand. Psychological bulletin 94, 2 (1983), 293.
@@
- Jung (2014) Carl Gustav Jung. 2014. Two essays on analytical psychology. Routledge.
@@
- Singh et al. (2024) Aneesha Singh, Marusa Hrobat, Suxin Gui, Nadia Bianchi-Berthouze, Judith Ley-Flores, Frederic Bevilacqua, Joaquin R Diaz Duran, Elena MÁrquez Segura, and Ana Tajadura-Jiménez. 2024. Pushed by Sound: Effects of Sound and Movement Direction on Body Perception, Experience Quality, and Exercise Support. ACM Transactions on Computer-Human Interaction 31, 4 (2024), 1–36.